OpenStudy (anonymous):

[(5 - 9) ∙ 3] (mod 5)

OpenStudy (anonymous):

first do the math for (5-9)3

OpenStudy (anonymous):

5 - 9 = ??

OpenStudy (anonymous):

help

OpenStudy (anonymous):

5 - 9 = ??

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

help

OpenStudy (anonymous):

what is a mod

OpenStudy (amistre64):

a mod is a remainder .....

OpenStudy (amistre64):

essentially; you want to know what remainer is of: [(5 - 9) ∙ 3] /5

OpenStudy (anonymous):

5-9=-4 -4x3=-12

OpenStudy (amistre64):

and -12/5 gives us a whole number and a remainder

OpenStudy (anonymous):

how do i do that

OpenStudy (amistre64):

you divide .... longhand style like back in the 4th grade

OpenStudy (amistre64):

since 5*-2 = -10, lets go for a smaller value and use 5*-3 = -15 -3 ----------- 5 ) -12 -(-15) ------ 3 <-- our remainder

OpenStudy (anonymous):

so r2

OpenStudy (amistre64):

the negative can be offputting; we want a value of 5*n such that it is less than -12

OpenStudy (anonymous):

but how did you get -15

OpenStudy (amistre64):

another trick is to count backlwards with a proper setup place your remainders in a row 0 1 2 3 4 now count backwards to -12 -12 -11 -10 -9 -8 -7 -6 -5 -4 -3 -2 -1 0 1 2 3 4 <-- poosible remianders of dividing by 5 ^^ -12 falls int his stack

OpenStudy (amistre64):

5*-1 = -5 ; -5 is bigger than -12 5*-2 = -10 ; -10 is bigger than -12 5*-3 = -15; -15 is smaller than -12 so use -3

OpenStudy (amistre64):

we cant get a remainder if we are trying to subtract a large from a small; which is why you have to be cognizant of the negative

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

if we have 12 mod 5, can you tell me what they would have been?

OpenStudy (anonymous):

i dont understand

OpenStudy (amistre64):

we just went thru how we defined mods ...

OpenStudy (amistre64):

we can look at it this way as well 12 = 5n + R 12 = 5(2) + R 12 = 10 + 2 12 mod 5 = 2 since we get a Remainder of 2 ---------------------------- -12 mod 5: -12 = 5n + R -12 = 5(-3) + R -12 = -15 + 3 -12 mod 5 = 3 since we get a Remainder of 3

OpenStudy (anonymous):

(11 x 4. + 6] (mod 8) so i 11x4=44 44+6=50 r is 6

OpenStudy (amistre64):

the 50 is good; lets write it out now as: 50 mod 8 50 = 8n + R well 8(6) = 48 8(7) = 56 .... too big, lets use the 6 50 = 8(6) + R 50 = 48 + R R has to equal 2 50 mod 8 = 2

OpenStudy (anonymous):

o so 8x7=56 but 6 is to much so i do 8x6=48 but it is 2 to small so the R is 2 2+48=50 50=50

OpenStudy (amistre64):

i can understand that thought process, yes :)

OpenStudy (anonymous):

cool i understand mods

OpenStudy (anonymous):

but what is a (clock 7)

OpenStudy (anonymous):

2 + (5 - 2) (clock 7)

OpenStudy (anonymous):

5-2=3 2+2=4

OpenStudy (amistre64):

clock arithemetic is another name for mods; normally a clock has 12 hours: clock 12 would then be mod12 clock 7 is just another expression for mod7 http://www-math.ucdenver.edu/~wcherowi/clockar.html

OpenStudy (amistre64):

2 + (5 - 2) (clock 7) 2 + (3) (clock 7) 5 (clock 7)

OpenStudy (amistre64):

5 = 7n + R 7(0) = 0 7(1) = 7 .. too big, go with 0 5 = 7(0) + R 5 = R

OpenStudy (anonymous):

thanks you are a "big help"

OpenStudy (amistre64):

youre welcome, im sure im going to go thru these in the Fall; gonna be taking number theory ....

OpenStudy (anonymous):

o well i bet you are good at it

OpenStudy (amistre64):

lol, i can be :) goood luck, i have to get